OpenStudy (anonymous):

WILL GIVE A MEDAL PLEASE HELP Which of the following is the equation of a circle whose center is at the origin and whose diameter has endpoints (-1, 2) and (1, -2)? x^2 + y^2 = 2 x^2 + y^2 = 3 x^2 + y^2 = 4 x^2 + y^2 = 5

OpenStudy (anonymous):

Note that equation of a circle is as such \[(x-h)^2+(y-k)^2=r^2\] where your centre is (h,k) so what will it be?

OpenStudy (anonymous):

It says the center is the origin meaning (0,0)

OpenStudy (anonymous):

I don't know, I'm so confused @iambatman

OpenStudy (anonymous):

You were given the end points, which will help you find the radius, which is 2 as the end points of the diameter are (-1,2) and (1,-2) which should give you a diameter of 4. You also know the center since it mentions center is at the origin meaning (0,0). \[(x-0)^2+(y-0)^2=2^2\]

OpenStudy (alekos):

I think the diameter is 2sqrt5